1. Consider the mechanical energy of a body in geostationary orbit above the Earth's equator, at rGS=42000 km.

Consider the mechanical energy of the same body on Earth at the South pole, at re = 6400 km. For this problem, we consider the Earth to be spherical. (Remember, the object at the equator is in orbit, the object at the Pole is not in orbit.)

G = 6.67×10^−11 Nm^2 kg^−2, and the mass of the Earth is M = 5.97×10^24 kg

What is the difference in the mechanical energy per kilogram between the two?

E = ___ MJ.kg^−1(to two significant figures, don't use scientific notation)
1
Expert's answer
2020-03-16T13:16:27-0400

Mechanical energy at the surface of the earth=


"\dfrac{GM}{R}=\dfrac{6.67×10^{−11} \times5.97×10^{24}}{(6400)\times10^3)}=\dfrac{39.8199\times10^{13}}{64\times10^5}"

"=\dfrac{39.82\times 10^8}{64}=6.22\times 10^7J/kg"


Mechanical energy per kg at height ="\\dfrac{GM}{(R+h)}\n\n\u200b\t\n + \n\\dfrac{\nv ^\n2}2\n \n\u200b"


"=\dfrac{GM}{(R+h)}+\dfrac{GM}{(R+h)}=\dfrac{2GM}{(R+h)}"


"=\dfrac{2\times6.67×10^{−11} \times5.97×10^{24}}{(6400+42000)\times10^3)}"


"=\dfrac{2\times 39.8199\times 10^{13}}{(462\times 10^5)}"


"=\dfrac{79.6398\times10^{13}}{462\times 10^5}=0.172\times 10^8 J/kg"


"=1.72\times 10^7 J/kg"


Hence the required difference="6.22\u00d710 \n^7\n \u22121.72\u00d710 \n^7\n =4.5\u00d710 \n^7\n J\/kg"